Janet Ludlow’s firm requires all its analysts to use a two-stage DDM and the CAPM to value stocks. Using these measures, Ludlow has valued QuickBrush Company at $63 per share. She now must value SmileWhite Corporation.

a. Calculate the required rate of return for SmileWhite using the information in the following table:

                                                                            December 2010

Quick Brush SmileWhite
Beta 1.35 1.2
Market Price $45.00 $32
Intrinsic Value $63.00 ?

Note: Risk-free rate = 4.5%; expected market return = 16%.

Instruction: enter your answer as a percentage rounded to 1 decimal place.

Required rate of return 18.3%

b. Ludlow estimates the following EPS and dividend growth rate for SmileWhite:

First three years: 16% per year
Years thereafter: 10% per year

Estimate the intrinsic value of SmileWhite in December 2010 using the table above and the two-stage DDM. Dividends per share in 2010 were $1.

Instruction: enter your answer as a decimal number rounded to 2 decimal places.

Year       Dividends

2010       $1.00

2011       $1.16

2012       $1.35

2013       $1.56

2014       $1.72

Intrinsic stock value in 2013:     $  .

Intrinsic stock value in 2010:     $  
Answer #1

a). According to the CAPM,

Required Return = Risk-free Rate + [Beta * (Expected Market Return - Risk-free Rate)]

= 4.5% + [1.2 * (16% - 4.5%)]

= 4.5% + [1.2 * 11.5%] = 4.5% + 13.8% = 18.3%

b-1). Intrinsic Stock Value in 2013(V2013) = D(2014) / (r - g)

= [$1 * (1 + 0.16)3 * (1 + 0.10)} / [0.183 - 0.10] = $1.72 / 0.083 = $20.69

b-2). Intrinsic Stock Value in 2010(V2010) = [D(2011) / (1 + r)] + [D(2012) / (1 + r)2] + [{D(2013) + V(2013)} / (1 + r)3]

= [{$1 * (1 + 0.16)} / (1 + 0.183)] + [{$1 * (1 + 0.16)2} / (1 + 0.183)2] + [$20.69 + {$1 * (1 + 0.16)3} / (1 + 0.183)3]

= $0.98 + $0.96 + $13.44 = $15.38